# Posterior Ramifications of Prior Dependence Structures

### 11 Dec 2023 v1

@@ 4.2 Practical Implications
Our simulations demonstrate the value in considering the copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$ when choosing between two prior dependence structures. In particular, we may want to consider the local optima for the copula densities. This numerical study also suggests that copulas have limited ability to reliably prompt more accurate and concentrated posterior distributions around a fixed parameter value $\boldsymbol{\theta}_{0}$. The true value of $\boldsymbol{u}_{0}$ is unknown in practice. And even if $\boldsymbol{u}_{0}$ is such that $c_{2}(\boldsymbol{u}_{0})>c_{1}(\boldsymbol{u}_{0})$, it does not guarantee that $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ will be closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ – for small or large sample sizes.
@@
Michimae and Emura’s (2022) numerical studies considered several levels of multicollinearity between the regression covariates for a single $\boldsymbol{\theta}_{0}$ value. They also considered relatively small sample sizes ranging from $n=20$ to 200. Our numerical results suggest that it is pertinent to consider a variety of $\boldsymbol{\theta}_{0}$ values and sample sizes $n$ before making general statements about the impact of the prior copula on the posterior distribution. We acknowledge that Michimae and Emura (2022) considered a more complicated model, which leveraged a hierarchical framework to specify the prior copula. Although not the focus of this article, Monte Carlo simulation could likely be used to explore the local optima of such prior copula density functions and extend the results from Theorem 2 to a hierarchical framework. This extension is one of several possible generalizations that could be made to how the analysis prior in (4) is defined.
@@ 5 Discussion
In this paper, we emphasized the importance of considering whether prior dependence structures can be retained a posteriori when assessing the adequacy of multivariate prior distributions. We proved a theorem and resulting corollary that help discard potential prior dependence structures for a parameter $\boldsymbol{\theta}$ that will be routinely or chronically rejected as data are observed in general settings. Discarding such dependence structures simplifies the prior specification process. This corollary examined whether the prior dependence structure for $\boldsymbol{\theta}$ is compatible with that induced by the likelihood function. With this corollary, we critiqued past recommendations that suggested flexible copula-based priors were suitable as analysis priors. We argued that these priors are better suited for design purposes. This paper emphasized copula-based priors, but the notion of chronic rejection is still applicable to multivariate prior distributions that are not explicitly defined using copulas, such as multivariate normal priors. Because correctly specifying the dependence structure a priori for an arbitrary model may be difficult, we examined how the choice of prior copula impacts the posterior distribution. We proved several results regarding how this choice of copula impacts the convergence of the posterior mode to a fixed parameter value $\boldsymbol{\theta}_{0}$. While examining the local optima of candidate copula density functions is valuable, our numerical studies showed that prior copulas should not be selected to improve the posterior’s ability to recover $\boldsymbol{\theta}_{0}$.
Future work could extend the results from this paper to hierarchical settings, in which the hyperparameters of the prior copula are themselves random variables. Moreover, copula-based prior specification could also be applied in more complex designs. For example, much of the existing literature on prior elicitation focuses on the single model case. However, many studies seek to compare data from two groups, where the data from group $k$ is generated from the model $m_{k}(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{k})$ for $k=1,2$. In the absence of information about one of the two groups, prior elicitation is an arduous task. When less is known about group 2, copulas might be used in conjunction with the prior information about group 1 to specify a joint prior for $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$. Such methods could meaningfully facilitate design prior specification. Our results from this paper could be extended to provide guidance on whether these prior distributions would also be suitable analysis priors.
@@
Section SM1 of the supplementary materials proves that the Fisher information for the conditional parameterization of the multinomial model is diagonal. The code for the numerical studies is provided on Github: https://github.com/lmhagar/PosteriorRamifications

### 23 Mar 2024 v2

@@ 4.3 Connections to Previous Contributions
Our simulations demonstrate the value in considering the copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$ when choosing between two prior dependence structures. In particular, we may want to consider the local optima for the copula densities. This numerical study also suggests that copulas have limited ability to reliably prompt more accurate and concentrated posterior distributions around a fixed parameter value $\boldsymbol{\theta}_{0}$. The true value of $\boldsymbol{u}_{0}$ is unknown in practice. And even if $\boldsymbol{u}_{0}$ is such that $c_{2}(\boldsymbol{u}_{0})>c_{1}(\boldsymbol{u}_{0})$, it does not guarantee that $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ will be closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for small or large sample sizes. Regardless of whether the prior dependence structure is a chronically rejected one, choosing a prior copula to improve posterior concentration does not appear to be a sensible objective.
We now contrast our conclusions with Michimae and Emura’s (2022) recommendations. Their numerical studies considered several levels of multicollinearity between the regression covariates for a single $\boldsymbol{\theta}_{0}$ value. For their three regression coefficients, the corresponding $\boldsymbol{u}_{0}$ value lies directly on the upper Fréchet-Hoeffding bound. Their $\boldsymbol{u}_{0}$ value does not necessarily correspond to a local maximum of the Clayton or Gumbel copula density functions they considered. However, the Clayton and Gumbel families of copulas accommodate positive prior dependence and their density functions are generally largest near the upper Fréchet-Hoeffding bound. Their single value is therefore similar to that used in case 2 of our study. They also considered relatively small sample sizes ranging from $n=20$ to 200. Given the results for case 2 in Figure 5, it is not surprising that their normal posterior medians for the regression coefficients were generally closer to their $\boldsymbol{\theta}_{0}$ values when using Clayton and Gumbel copulas. Our numerical results suggest that it is pertinent to consider a variety of $\boldsymbol{\theta}_{0}$ values and sample sizes $n$ before making general statements about the impact of the prior copula on the posterior distribution.
@@
We acknowledge that Michimae and Emura (2022) considered a more complicated model, which leveraged a hierarchical framework to specify the prior copula. Although not the focus of this article, Monte Carlo simulation could likely be used to explore the local optima of such prior copula density functions and extend the results from Theorem 2 to a hierarchical framework. This extension is one of several possible generalizations that could be made to how the analysis prior in (4) is defined.
@@ 5 Discussion
In this paper, we proposed a framework to consider whether prior dependence structures can be retained a posteriori. This framework improves transparency when making inference about posterior dependence structures and helps discard chronically rejected dependence structures for a parameter $\boldsymbol{\theta}$ that cannot be retained as data are observed. Discarding such dependence structures simplifies the prior specification process, particularly when practitioners aim to collect large samples. We also discussed small-sample objectives for prior specification to clarify whether the inability to retain the prior dependence structure presents practical issues for a given posterior analysis. This paper emphasized copula-based priors, but the notion of chronically rejected dependence structures is still applicable to multivariate prior distributions that are not explicitly defined using copulas, such as multivariate normal priors.
Since correctly specifying the dependence structure a priori for an arbitrary model may be difficult, we examined how the choice of prior copula impacts the posterior distribution. We proved asymptotic results regarding how this choice of copula impacts the convergence of the posterior mode to a fixed parameter value $\boldsymbol{\theta}_{0}$. While examining the local optima of candidate copula density functions is valuable, our numerical studies showed that prior copulas should not be selected to improve the posterior’s ability to recover $\boldsymbol{\theta}_{0}$. These results contradicted past recommendations that suggested the choice of prior copula could reliably improve posterior concentration.
Future work could extend the results from this paper to hierarchical settings, in which the hyperparameters of the prior copula are themselves random variables. Moreover, the theoretical results in this paper are based on the standard Bernstein-von Mises theorem. To broaden the applicability of this framework, we could consider nonparametric methods for specifying prior dependence. We could also consider chronically rejected prior dependence structures in the presence of model misspecification (i.e., when $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y})$ that defines the posterior does not coincide with $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ used to generate the data). In that case, we require more complex characterizations of prior dependence than $\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}$ for large samples.
@@
Appendix A proves the Fisher information for the conditional parameterization of multinomial model is diagonal. Appendix B contains additional simulations for the calibration of credible sets. Appendix C proves Theorem 2. The code for the numerical studies is provided on Github: https://github.com/lmhagar/PosteriorRamifications